# version 0.75

### v4

@@ 13.1 The Gradient Response Controller
Now, we use the fact that the iterates of OGD $M_{1:h}^{t}$ move slowly to establish that $\mathbf{y}_{t}$ and $\mathbf{y}_{t}(M^{t}_{0:h})$ are close in value. Before that, by definition, we have $\mathbf{u}_{t}(M^{t}_{0:h})=\mathbf{u}_{t}$ since the sequence $\mathbf{y}^{\mathbf{nat}}_{t}$ is determined independently of the choice of the policy being executed. Due to the $\gamma$-stability of the underlying system, we have for any $i>0$, $\|C{A}^{i}B\|\leq\kappa(1-\delta)^{i}$ for some $\kappa>0,\gamma<1$. Using Lemma 12.7, we have for some $C>0$ that
@@
The control depends only on the parameters of the current iteration and we have that $\mathbf{u}_{t}=\mathbf{u}_{t}(M_{1:h}^{t})$. By definition, $l_{t}(M_{0:h}^{t})=c_{t}(\mathbf{y}_{t}(M_{0:h}^{t}),\mathbf{u}_{t}(M_{0:h}^{t}))$. Thus, we have
@@ 13.2 Extension to Time Varying Systems
The GRC algorithm can be extended to time-varying dynamics as we explain in this section. The linear dynamical system at time $t$ is denoted by system matrices $A_{t},B_{t},C_{t}$, and we assume that they are known to the controller. We further assume that these dynamical systems are all $\gamma$-stable. The algorithm is formally spelled out in figure 9.
@@ 13.3 Conclusion
The results in this chapter demonstrate that adaptive control under partial observability is possible without explicit state estimation, bridging the gap between classical observer-based controllers and modern learning-based methods. By leveraging regret minimization techniques, we provide a robust and flexible approach that applies even in adversarial and uncertain environments.
@@ 13.4 Bibliographic Remarks
The setting was further extended to changing linear dynamics in [MGSH21], where the expressive power of different policy classes was also studied.
The analogue of online nonstochastic control in the stochastic noise setting is called the Linear Quadraic Gaussian problem, and it is further limited by assuming quadratic noise. This classical problem in optimal control was proposed in Kalman’s work [Kal60]. Regret bounds for controlling unknown partially observed systems were studied in [LAHA20b, LAHA20c, MTR19]. [LAHA20a], which uses DRC policies even in context of stochastic control, obtains rate-optimal logarithmic bounds for the LQG problem.
Control with partial information and bandit feedback with semi-adversarial noise was considered in [SNH23]. Tight rates for quadratic control with bandit feedback, in the fully adversarial model, were obtained in [SSNH24], and more general cost functions in the bandit setting were considered in [SL24].
@@ 13.5 Exercises
13.1. In this exercise we consider the computation of nature’s y’s for time varying systems. Let
@@
13.3. Prove that the optimal controller that minimizes the aggregate cost for a partially observed linear dynamical system assuming the perturbations are stochastic is a linear dynamical controller.
13.4. Prove that the regret of the GRC controller for time varying linear dynamical systems is bounded by $O(\sqrt{T})$.

### v5

@@ 13.1.1 Finishing up the Regret Bound
Now, we use the fact that the iterates of OGD $M_{1:h}^{t}$ move slowly to establish that $\mathbf{y}_{t}$ and $\mathbf{y}_{t}(M^{t}_{0:h})$ are close in value. Before that, by definition, we have $\mathbf{u}_{t}(M^{t}_{0:h})=\mathbf{u}_{t}$ since the sequence $\mathbf{y}^{\mathbf{nat}}_{t}$ is determined independently of the choice of the policy that is being executed. Due to the $\gamma$-stability of the underlying system, for any $i>0$, $\|C{A}^{i}B\|\leq\kappa(1-\delta)^{i}$ for some $\kappa>0,\gamma<1$. Using Lemma 12.7, we have for some $C>0$ that
@@
The control depends only on the parameters of the current iteration, and we have $\mathbf{u}_{t}=\mathbf{u}_{t}(M_{1:h}^{t})$. By definition, $l_{t}(M_{0:h}^{t})=c_{t}(\mathbf{y}_{t}(M_{0:h}^{t}),\mathbf{u}_{t}(M_{0:h}^{t}))$. Thus, we have
@@ 13.2 Extension to Time Varying Systems
The GRC algorithm can be extended to time-varying dynamics as we explain in this section. The linear dynamical system at time $t$ is denoted by the system matrices $A_{t},B_{t},C_{t}$, and we assume that they are known to the controller. We further assume that these dynamical systems are all $\gamma$-stable. The algorithm is formally spelled out in figure 9.
@@ 13.3 Conclusion
The results in this chapter demonstrate that adaptive control under partial observability is possible without explicit state estimation, bridging the gap between classical observer-based controllers and modern learning-based methods. Using regret minimization techniques, we provide a robust and flexible approach that applies even in adversarial and uncertain environments.
@@ 13.4 Bibliographic Remarks
The setting was further extended to time-varying linear dynamics in [MGSH21], where the expressive power of different classes of policy was also studied.
The analogue of online nonstochastic control in the stochastic noise setting is called the Linear Quadraic Gaussian problem, and it is further limited by assuming quadratic noise. This classical problem in optimal control was proposed in Kalman’s work [Kal60]. Regret bounds for controlling unknown partially-observed systems were studied in [LAHA20b, LAHA20c, MTR19]. [LAHA20a], which use DRC policies even in the context of stochastic control, obtain rate-optimal logarithmic bounds for the LQG problem.
In [SNH23], the task of controlling a partially-observed system subject to semi-adversarial noise under bandit feedback was considered. The tight rates for quadratic control with bandit feedback, in the fully adversarial model, were obtained in [SSNH24], and more general cost functions in the bandit setting were considered in [SL24].
@@ 13.5 Exercises
13.1. In this exercise, we consider the computation of nature’s y’s for time-varying systems. Let
@@
13.3. Prove that the optimal controller that minimizes the aggregate cost for a partially observed linear dynamical system assuming that the perturbations are stochastic is a linear dynamical controller.
13.4. Prove that the regret of the GRC controller for time-varying linear dynamical systems is bounded by $O(\sqrt{T})$.